myabell:

Jonathan wants to buy a lid for his frying pan. The diameter of the pan is 9 inches. What is the approximate circumference of a lid that will fit his frying pan? (Use 3.14 for .)

Extrinix:

Do you know the equation for the circumference of a circle?

myabell:

I think it is 9 x 3.14

myabell:

or 9 divides 3.14

Extrinix:

\(2\pi r^2\) or \(\pi d\) So you would be correct yeah.

Extrinix:

So what would the answer of \(2\pi (4.5)^2\) or \(\pi (9)\) be?

Extrinix:

Correction* \(2\pi r\). (Copy and paste be like.)